### v1

# Semantic Inference-Based Deep Learning and Modeling for Earth Observation: Cognitive Semantic Augmentation Satellite Networks

Abstract
Earth Observation (EO) systems support vital tasks such as mapping, disaster monitoring, and resource management, yet face challenges in effectively processing and transmitting vast volumes of EO data, particularly in specialized applications like agriculture and real-time disaster response. Domain-adapted Large Language Models (LLMs) offer a solution by enabling data fusion between mass EO data and semantic EO data. Through domain adaptation, LLMs improve the integration and interpretation of diverse data sources, addressing the challenge of processing specialized datasets in applications like agriculture and disaster response. The fusion of these data types, guided by LLMs, enhances the accuracy and relevance of the transmitted information. This paper presents a comprehensive overview of leveraging semantic inference for deep learning and semantic modeling for advanced EO systems. A novel framework is presented for semantic communication in EO satellite networks, aimed at enhancing data transmission efficiency and system performance through cognitive processing techniques. The proposed system leverages Discrete Task-Oriented Joint Source-Channel Coding (DT-JSCC) and Semantic Data Augmentation (SA) integrate cognitive semantic processing with inter-satellite links, enabling efficient analysis and transmission of multispectral imagery for improved object detection, pattern recognition, and real-time decision-making. Cognitive Semantic Augmentation (CSA) is introduced to enhance a system’s capability to process and transmit semantic information, improving feature prioritization, consistency, and adaptation to changing environments and application needs. The end-to-end architecture is designed for next-generation satellite networks, such as those supporting 6G, demonstrating significant improvements in efficiency and accuracy.
Index Terms:
Earth observation, remote sensing, semantic data augmentation, on-board processing, deep joint source-channel coding, cognitive semantic satellite networks, semantic communication, large language model
I Introduction
The United Nations has acknowledged Earth observation (EO) as a crucial instrument for accomplishing many Sustainable Development Goals [1]. EO systems acquire and analyze data about the globe for numerous civil applications and critical tasks that significantly impact human existence through the use of satellite networks. These tasks include mapping, weather forecasting, risk management, disaster monitoring, natural resource management, and emergency response. The estimated value of the satellite-based EO market is expected to reach 11.3 billion USD by 2031 [2]. The Internet of Things (IoT) has led to significant advancements in information and communication technology, which has increased demand for profitable value-added services and applications. The emergence of big EO data has been facilitated by the progress in sensors, communication, and storage technology. This dataset has a wealth of valuable information, but it also presents some difficulties owing to its distinctive characteristics, such as one-time observations and intricate dimensions. The exploration of picture retrieval strategies is tailored to address the challenges posed by complicated data [3]. The study emphasizes notable obstacles, including the scarcity and diversity of training data, as well as the current algorithms’ inability to imitate human-like cognition.
Although there is a large amount of EO data available, there is a lack of effective conversion into actionable information for specialized sectors such as agriculture or disaster response. Big data systems that have frequent satellite revisits have the potential to revolutionize real-time monitoring [4]. NewSpace constellations [5] are expanding their capabilities beyond the provision of affordable and frequent data. They focus on addressing particular requirements by using specialized satellites and providing data analysis in addition to raw data, serving a broader spectrum of consumers. Furthermore, monitoring strategies need careful consideration of both geographical and temporal dimensions [6]. In terms of space, this entails establishing the specific region or extent that is of interest. The extent of the phenomena being examined might vary from a single spot to an intricate three-dimensional grid. Resolution, which refers to the level of detail in measurements, and expanse, which indicates the region covered, are both important aspects. An example of researching the influence of topography on soil moisture may include the measurement of moisture levels at many locations along a slope, necessitating a high level of spatial resolution. Consequently, current EO systems are facing difficulties in managing the substantial volume of data being sent and fulfilling the diverse demands for the latency at which the data has to be processed. In order to address these significant challenges, the primary strategies involve two actions: firstly, improving the bandwidth of satellite communication networks that facilitate EO systems, and secondly, reducing data volume by employing on-board processors on the satellites to preprocess the EO data.
Utilizing the inter-satellite linkages (ISLs) of the multi-satellite, multi-orbit connectivity systems has been seen as a potential strategy, as stated in the first direction [7]. The integration of several satellites and orbits, spearheaded by non-geostationary satellites (NGSO), is on the verge of transforming worldwide communication. Unlike conventional geostationary satellites (GSO), these systems provide reduced latency and signal loss as a result of their lower orbits. These benefits include the provision of fast internet access in isolated regions, improved capacity for transmitting data via 5G networks [8], and the possibility of establishing direct, high-capacity connections with consumers. Nevertheless, the smooth integration with current GSO and terrestrial networks requires careful and detailed design due to several constraints. Efficient cohabitation across systems needs solutions for spectrum access and restrictions [9]. An additional challenge lies in the efficient design and management of extensive constellations of NGSO [10]. It is crucial to build user equipment that can successfully establish communication with these constellations. This research examines the communication in multi-satellite, multi-orbit connectivity systems, with a special emphasis on NGSO technology. It investigates the possibilities of this technology while recognizing the obstacles associated with its adoption. The data acquired by EO satellites may undergo processing, storage, and transmission to Earth using SatCom satellites positioned in various orbits. In the context of this framework, a remote sensing satellite (or EO satellite) refers to a spacecraft that collects EO-related data from distant sensors and IoT devices. By using the ISLs, several routing channels may be constructed between the remote-sensing satellites and their respective ground stations. The technique of [11] may effectively handle the problem of overload and backhaul bottlenecks that occur at certain nodes in inter-satellite communication (ISC) networks, resulting in improved network capacity.
The transmission of the meaning behind data is the primary objective of semantic communication, rather than the precise data itself. This has the potential to enhance efficiency by providing greater flexibility in encoding and decoding. The authors of [12] suggest a method that is predicated on three concepts: the use of a specialized data format to capture meaning, the adaptation of transmission based on channel conditions, and the use of generative models to adjust transmission and enable tasks at the receiver. Researchers of [13] have devised a system that utilizes natural language conversations to disseminate flood risk information, with the aim of enhancing public preparation for floods. This system utilizes a large language model (LLM) that has been trained specifically on flood-related knowledge, guaranteeing the precision and reliability of the information it provides. Through the use of geographic information systems (GIS), the LLM has the ability to tailor risk evaluations according to the geographical location of the user. Utilizing natural language conversations has shown the potential to enhance flood risk perception among individuals with diverse cognitive capacities. Furthermore, semantic Importance-Aware Communication (SIAC) uses LLM to understand the meaning and semantics of conveyed data [14]. This understanding allows the system to give priority to significant information by allocating more resources to data with greater semantic significance. Unlike earlier approaches, SIAC can seamlessly integrate into current communication networks with minimal alterations. By prioritizing crucial information, SIAC reduces semantic loss more effectively than conventional communication techniques.
In relation to the second aspect, the use of semantic communication (SemCom) principles has been recognized as a favorable strategy. The primary purpose of SemCom is to prioritize semantic accuracy above reducing bit/symbol error rates at the receiver side, which is in complete alignment with the core goals of most EO methods. The first principle of SemCom is the extraction of semantic data, wherein pertinent information is initially obtained, condensed, or broken down from unprocessed data. Afterward, just the information related to meaning is transferred to the central core. The PhiSat-1 project presented in [15] by the European Space Agency (ESA) might be regarded as the first demonstration of this idea. Two 6U CubeSats are deployed in this mission to capture Earth photographs. The data is processed inside the CubeSats to filter out information depending on the cloud coverage of the scene. The second concept revolves around semantic-oriented network management, specifically SemCom. SemCom’s objective is to oversee all network functionalities across various layers (PHY, MAC, Network, Transportation, Application) by taking into account the semantics of the data and the requirements of applications related to EO tasks [16].
Irrespective of the intended message of these bits, the aspiration of beyond Shannon [17] beyond the traditional Shannon paradigm, which emphasizes the reliable reception of individual sent bits. Semantic-information feature learning [18] is the key to addressing the utilization of cognitive techniques employed as a means to direct computational resources towards activities of higher importance. The process of achieving dynamic adaptation in semantic compression involves the utilization of a feature learning module and an attention feature module. These modules enable the source encoder to generate a variable number of symbols and modify the capabilities of the source encoder and channel encoder through the use of cognitive techniques. The evolution of semantic communication [19] is traced back to the early 20th century with its continuous growth into the realm of modern communications beyond 5G and 6G technologies. In light of this, there is a significant need to develop more intelligent and efficient communication protocols that can meet the diverse quality of service (QoS) needs. This must be done while addressing the challenge of limited communication bandwidth and computation. The development of an intelligent communication system is considered essential in both industry and academia. Such a system is not limited to memorizing data flows based on rigorous regulations but also aims to comprehend, analyze, and articulate the fundamental semantics. In order to facilitate the implementation of immediate intelligent services in future communication systems, it is advantageous to distill and communicate only the information that is pertinent to the task at hand and precise semantics. The semantic approach effectively reduces the overall latency of the system.
Figure 1: The system model of semantic inference-based earth observation satellite networks.
Fig. 1 presents a conceptual model for a multi-orbit inter-satellite network in the context of Earth observation, highlighting a layered approach to satellite communication and data processing. At the center, Earth observation satellites, operating in different orbits (LEO, MEO, and GEO), gather raw data through ground sensors distributed across various landscapes, such as agricultural fields and maritime regions. This raw data is transmitted to a ”Ground segment,” which includes data centers and user control interfaces for further processing. Data travels through multiple layers of communication within the network. At the base level, the PHY Layer (Physical Layer) ensures the fundamental transmission of data, followed by the MAC Layer (Medium Access Control) responsible for resource allocation. Moving upwards, the Network Layer manages the interconnectivity across the satellite network. The Transportation Layer handles the movement of data across the system, connecting the ground segments to various orbital satellites.
Above these operational layers lies the Semantic Effectiveness Plane, which provides a high-level oversight and analysis of the semantic data flow. Two processes stand out at either end of this system: Semantic Extraction and Semantic Restoration. Data extracted at the ground level undergoes semantic processing, where meaningful patterns or insights are identified and interpreted. This processed semantic information is later restored into useful applications, ranging from environmental monitoring to disaster management. The model suggests a closed-loop system where raw data from ground sensors is transformed into semantically rich insights, enabling diverse applications through seamless communication between satellite networks and ground infrastructure.
We summarize the paper as follows. Section I discusses semantic image processing, focusing on pixel-level categorization to improve applications like autonomous driving, medical imaging, and Earth observation while addressing challenges with complex data and deep-learning techniques for more accurate analysis. Section II examines semantic communication networks and their integration into 6G systems, emphasizing goal-oriented methods and joint source-channel coding to enhance communication efficiency. Section III explores adapting LLMs for domain-specific tasks to optimize semantic transmission, while Section VI presents the Cognitive Semantic EO system model, integrating DT-JSCC and semantic data augmentation for efficient EO data processing in satellite networks.
II Semantic data processing for Earth observation and 6G communication
This section delves into the concept of semantic segmentation, a technique that assigns precise identifiers to individual pixels in photographs. It concentrates on the application of this methodology to the examination of satellite imagery and other types of EO data. The content explores the application of semantic segmentation in a variety of applications, including trajectory facilitation, information retrieval, and picture compression. We also underscore the current advancements in deep learning for the classification of various EO data types using semantic knowledge.
Assigning a unique name to each pixel in an image is a basic task in computer vision known as semantic image processing. Achieving accurate categorization at the pixel level is crucial in applications like driverless cars, where the capacity to understand lane markings and precisely identify the position of objects is essential. However, the process of semantic segmentation may be complicated by barriers such as hidden objects or congested backgrounds. Notwithstanding these challenges, semantic segmentation offers benefits that go beyond autonomous vehicles. The technology enables the examination of medical images by identifying abnormalities, assists robots in navigating their surroundings, and improves video surveillance by detecting potentially troublesome behavior. Recent progress in deep learning has led to remarkable achievements in achieving precise precision at the pixel level. However, its performance is inferior to that of normal satellite images when dealing with non-standard data such as point clouds and spectrum imaging. Likewise, a significant disparity in performance occurs when dealing with limited data sets. The lack of comprehensively classified, unconventional remote sensing data hinders the development and evaluation of innovative deep-learning methods.
II-1 Computational semantic compression
The study presented in [20] investigates potential resolutions for constraints such as streamlining features, indexing, semantic labeling, and ultimately, approaches for formulating and executing searches. As stated in [21], the primary objective of the study is to evaluate the existing condition of research on extracting information from EO data. The authors of [22] provide a new notion of a semantic EO data cube. It enriches traditional exploratory optimization data cubes by including an extra layer, referred to as a ”nominal” or categorical interpretation, for each observation. Thus, it becomes feasible to use semantic searches that directly relate to the substance of the data and are easily comprehensible to people. By using machine-based reasoning, this approach improves automated analysis and information extraction in comparison to non-semantic cubes. Fundamentally, semantic EO data cubes enable a wider range of users to utilize comprehensive EO data. This study evaluates CORSA [23] and [24], an innovative compression method powered by artificial intelligence that produces useful data representations at the same time. The authors compare traditional methods for detecting water with innovative ways that use the compressed data and representations of CORSA. The results suggest that CORSA compression does not have any detrimental effect on accuracy and allows for the direct generation of subsequent water detection networks based on its representations. These results demonstrate the adaptability of CORSA as a solution for processing and compressing data on-board, successfully addressing the growing challenge of managing EO data. Furthermore, the authors of [25] propose a novel method that uses semantic technology to provide a standardized language for EO data. The methodology emphasizes the smooth integration of data obtained from raster imagery, such as land cover maps, with other geographical information. This enhances the potential for more efficient analysis and use of this crucial EO data.
The exponential growth of data produced by satellites and other remote sensing devices poses a new challenge: the rise of vast Earth observation data. Despite the great potential of this data for many applications, such as the provision of disaster assistance, its processing and interpretation present substantial difficulties. The study presented in [3] investigates existing methods for categorizing distinct images from this extensive collection, highlighting their practical usefulness. The authors examine potential methodologies for analyzing and managing large-scale EO data by using advancements in computer vision, machine learning, and other disciplinary domains. The study described in [26] investigates the many factors that might influence the effectiveness of deep features for image retrieval. By optimally adjusting these factors, the authors of [26] may get exceptional results on publicly available datasets. The present work provides valuable insights for future efforts in retrieving high-resolution remote sensing (HRRS) photographs by analyzing their content.
II-2 Fusion of semantic trajectories
Global Positioning System (GPS) data produces huge amounts of trajectory information. While scholars have largely focused on the preservation and analysis of these trajectories, there is a growing need to understand the underlying importance of the movement. The material includes traditional compression algorithms, the criteria for defining moving objects, approaches for measuring compression efficiency, and potential future applications. For real-time applications, conventional methods for producing substantial movement data are inadequately fast. The research presented in [27] presents a novel approach that combines mobility data with geography and application information to produce semantic trajectories. These improved trajectories not only reveal the destinations of objects but also provide a valuable understanding of the fundamental causes of their motion.
The SeTraStream system, as described in [28], tackles this problem by analyzing real-time semantic trajectories generated from live streaming movement data. This enables prompt processing for applications like location-based queries. Furthermore, the study shown in [29] used an automated method to classify points in trajectories that display similar characteristics, leading to a more concise and meaningful depiction. This methodology is adequately effective in handling live data streams, making it suitable for real-time applications. An expanding array of advanced technologies described in [30] and [31] have become available for collecting and combining an unprecedented amount of data from numerous sources, including GPS trajectories related to the routes of mobile objects. Thus, the COMPRESS system described in [32] is specifically developed to efficiently compress GPS data collected from urban road networks. The COMPRESS algorithm works by first partitioning the data into geographical routes and timestamps, then thereafter compressing each individual component independently. Similar to the method described in [33], this technique achieves significant compression ratios while allowing location-based searches on the compressed data. Nevertheless, the authors of [33] compress the data by identifying and keeping just the most important components based on these criteria. This methodology effectively reduces the amount of data while maintaining crucial information for further study of ship movement patterns.
II-3 Advancements in Semantic Comprehension
To optimize the comprehension of movement data (trajectories) by incorporating semantic understanding, two crucial methods to consider are compression, which aims to decrease the volume of data, and semantic segmentation, which involves dividing trajectories into meaningful segments annotated with semantic labels. This paper focuses on emerging developments, including the integration of additional data dimensions and the facilitation of real-time processing. The preferred approach is contingent upon many elements, including urgent needs and the equilibrium between reducing the volume of data and comprehending its importance. Furthermore, the work outlined in [34] presents a novel approach to integrating semantic information to enhance trajectory simplification. The approach does this by first distinguishing and delineating periods of movement from occurrences of halting within trajectories. Moreover, these sections are then simplified individually. The HighDAN AI network, as outlined in [35], is built around this approach and has the capability to comprehend the significance of content conveyed by intricate imagery, such as satellite images, originating from various urban areas. The system exhibits remarkable competence in managing the wide range of ways in which these images depict identical geographical features in diverse urban settings. These capabilities allow HighDAN to attain exceptional performance and generalizability in the process of semantic compression from remote sensing data.
The research conducted in [36] directly investigates the use of semantic segmentation in three-dimensional meshes. Semantic segmentation is the process of categorizing different sections of a mesh into certain groups, such as urban structures or roadways. While deep learning has shown efficacy in assessing many types of remote sensing data, such as point clouds and satellite photos, there is a lack of assessments particular to 3D mesh segmentation. The present evaluation aims to fill the existing research gap regarding the use of deep learning techniques for the purpose of semantic segmentation of extensive 3D models inside urban environments. The authors explore the application of deep learning approaches to meshes, different ways of classification, and the assessment of experimental results [36]. The SAMRS framework, as described in [37], is a novel system that includes a very large collection of over 100,000 photographs. The photos undergo automated recognition and include exact EO data of superior quality. The SAMRS system utilizes existing object detection data and applies the Segment Anything Model (SAM) approach [38] to effectively provide semantic labels and segmentation. This dataset has many benefits, including its vast size, comprehensive item information, and adaptability for incorporating various semantic segmentation tasks.
TABLE I: Summary of Methodologies and Applications in Semantic Image Processing and Compression
Feature [20] [21] [22] [23] [25] [3] [26] [27] [29] [30] [32] [34] [35] [36] [37] Proposed
EO data processing
AI-based compression
Semantic labeling
Trajectory analysis
Image retrieval
Data fusion
Geospatial analysis
Environmental monitoring
Disaster response
Remote sensing
Compression
On-board satellite processing
Movement prediction
3D semantic segmentation
Standardization
High-resolution sensing
Vessel movement data analysis
Urban mobility analysis
Maritime logistics
Cross-urban analysis
Satellite data compression
Deep feature optimization
Stream analysis
Mobility data analysis
GPS data processing
Key component extraction
Handling diverse urban data
Water detection networks
Categorization for disaster assistance
Table I highlights the significant advancements in semantic image processing, computational semantic compression, and the fusion of semantic trajectories within the context of EO and related fields. From streamlining data searches and enhancing data integration to innovative compression methods like CORSA, these contributions emphasize the growing importance of semantic technologies in handling large-scale EO datasets. The development of semantic EO data cubes, real-time semantic trajectory processing systems, and advanced AI models such as HighDAN and SAMRS demonstrates how semantic understanding is transforming data analysis, enabling more efficient and accurate results. These innovations pave the way for improved applications across various domains, from autonomous vehicles and urban planning to medical image analysis and video surveillance. Nonetheless, challenges remain, particularly with respect to non-standard data types and limited datasets, requiring further research to optimize deep learning models and enhance generalizability across diverse environments.
III Formation of semantic communication networks
The utilization of semantic communication (SemCom), which specifically prioritizes the transmission of the fundamental meaning (semantics) of information, has been increasingly popular because of its ability to create more efficient communication systems. Notwithstanding this commitment, there remains a discrepancy between theoretical principles and their practical execution. In [39], the authors present that fills the existing gap by reevaluating the core concepts of SemCom and their impact on system design. The investigation includes two primary components: a goal-oriented methodology that gives priority to metrics focused on data interpretation, and semantic operability, which considers both data sharing and its relevance. Within the field of engineering design, semantic networks play a crucial role by using knowledge graphs called nodes and links to visually represent ideas and their interconnectedness [40]. They are increasingly employed as storage places of information to support various engineering design tasks.
III-1 Semantic networks evolving toward SemCom networks
An inherent attribute of artificially intelligent systems intended for various applications is their capacity to premeditate goal-oriented actions based on predetermined conditions of a problem space. The development of transformation, substance, and interpretation operations for semantic networks is aimed at organizing the behavioral planning of self-governing intelligent systems during the resolution of problematic issues [41]. These procedures are supported by a formalization of the current situations in a problem space characterized by vast dimensions. Semantic networks aim to verify the understanding of a semantic representation and the ability to create or enhance a semantic language that corresponds to the viewpoint of the educational instructor. The engineering design field makes use of two main classifications of semantic networks:
  • WordNet [42] and ConceptNet [43] are representative instances of general knowledge networks that serve as repositories for common-sense knowledge. These networks have the potential to be very beneficial for tasks such as idea generation and information retrieval. Nevertheless, they may be deficient in the requisite level of comprehensiveness for some technological fields.
  • Engineering and technical networks, as documented in [44], such as B-Link and TechNet, are specifically designed for engineering applications and include technical expertise and professional networks. Their capacity to improve efficiency in activities such as patent search and design review is considerable.
The restricted availability of infrastructure resources is a significant challenge in meeting the extensive demand with stringent performance standards for the goal of semantic networks. SemCom works by using information bases, which are collections of background and contextual data that are shared ahead of time between machines communicating to arrive at a shared understanding of the communication goal [45]. The SemCom networks enable a wide connection without requiring a common vocabulary or structure for data transfer, therefore improving communication effectiveness and reliability, resulting in higher Quality of Service (QoS) especially tailored to individual requirements.
Following the SemCom networks presented in [46], a semantic-aware infrastructure provides provisioning and orchestration for 6G services with stringent QoS/QoE requirements. Service providers register their services, which consist of functional instances such as delivering and monitoring of motion, and this registration is facilitated via SemCom through relevant Knowledge Bases (KBs). These services use lifelong learning activities to ensure the knowledge repositories are regularly updated. Clients seek access to these services by establishing direct connections that manage the flow of semantic information to the service instances. The reasoning plane in end nodes in advanced language maturity requires semantic depictions from the user and control planes for the succeeding multiple access command. This enhances the cognitive capacities and operational autonomy of radio nodes in their decision-making processes for networking and control.
III-2 Orientation and attributes of semantic communication
In light of the limitations imposed by traditional methods, the SemCom systems of 6G need a more sophisticated approach to resource allocation, especially in dynamic environments. Conventional methods of communication sometimes overlook the importance of the fundamental relevancy of information. Their operation is based on the concept that each bit or symbol has equal significance and is considered as such. A key goal of these systems is to guarantee precise sequence retrieval at the receiving ends, with a focus on prioritizing the transmission of conformance. The predominant design techniques in this domain have been established based on ideas derived from digital communication. Theoretical limits of information determine the highest possible capacity of a certain system. The goal of channel coding is to devise techniques that can achieve these restrictions with a very low probability of errors. Conversely, source coding, often known as data compression, involves encoding data in a manner that reduces the necessary data size to maximize the length of the encoded sequence from the source. Nevertheless, the latest manifestation of communication systems is being executed in manners that question the traditional design paradigm, particularly with regard to semantic elements.
In [47], deep reinforcement learning is used to tackle the challenges encountered in multi-cell systems. By translating data rates into semantic information rates and utilizing a unique metric called Semantic Throughput (STP), the technique improves resource allocation and achieves improved STP performance compared to standard methodologies. This not only improves the effectiveness of communication, but also guarantees the dependable transmission of meaning, therefore laying the foundation for advancements in 6G. As Shannon’s Law nears its ultimate limits, researchers are exploring the theoretical constraints of compressing substantial information (semantic information). The Bayesian network used to represent the semantic source evaluates both lossless and lossy compression techniques [48]. The authors establish theoretical limits for attaining perfect reconstruction and propose an efficient encoding model that depends on the network’s structure. Moreover, they provide the most efficient compression methods for scenarios where information is available to both the sender and receiver. They illustrate that the optimal method entails incorporating distinct enhancements throughout the network. This work provides a basis for understanding the maximum degree of compression that may be successfully achieved for semantic data.
The semantic attribute assesses the congruence between the recipient’s deduced understanding and the presenter’s intended message, by considering the content, specifications, and semantics, with the aim of improving the intelligence of the communication system. Coherence encompasses both the precision and lucidity with which symbols may convey their intended meanings, as well as the effective conveyance of information or a concept from its origin to its destination, without delving into particular particulars. The evaluation of the compression level attained in semantic source coding in relation to the original information is of utmost significance [49]. Through the use of this semantic source, the authors have derived the theoretical limits for both lossless and lossy compression, along with the minimum and maximum constraints on the rate-distortion function. Based on the defined research objectives and existing knowledge, it can be shown that semantic source coding [19] has the potential to achieve more significant reductions in communication overhead and redundancy cost when compared to separate source-channel coding (SSCC). The primary factor for this is its ability to collect and compress the most pertinent and concise information, and then summarize it effectively.
  • Evaluation of the level of compression achieved in semantic source coding with respect to the original information is crucial. In order to determine the theoretical boundaries for lossless and lossy compression, the writers of [49] employed this semantic source. Moreover, they established the lower and upper limits for the rate-distortion function.
  • The aim of the task-oriented component is to effectively accomplish tasks while taking into account the possible influence on the value of the information delivered. Leveraging easily accessible resources such as communication bandwidth, processing expenses, and energy consumption enhances the feasibility of attaining a shared goal within the constraints and specifications of a particular project.
  • An assessment of the effectiveness of a task or performance indicator relies on the extent to which the obtained information contributes to its achievement. System performance evaluation may be quantified by measuring the extent to which a certain objective is achieved, taking into account the allocated resources.
These previously described perspectives differ from the approach grounded on the information theory paradigm, which considers all potential transmission sequences.
III-3 Semantic joint source-channel coding
Figure 2: The block diagram for semantic inference-based earth observation.
Traditional communication networks treat all data as equal and provide priority to sending it without any mistakes. This statement neglects to recognize the varying importance of the information itself. Source coding is a data compression technique used to prioritize the delivery of just the essential components. This suggests that conventional methods may be improved by giving priority to the relevancy of information. The objective of SemCom is to understand the fundamental meaning of the message. Despite its growth, the discipline has challenges such as developing models for complex data and using current communication systems. Notwithstanding the obstacles, the Joint Source-Channel Coding (JSCC) method has the capacity to enhance semantic communication in future systems like 6G.
The partnership between JSCC and deep learning, as presented in [50], showcases that the deep learning encoder and decoder outperform the traditional approach in terms of word error rate, particularly when there are limitations on the computational resources designated for syllable encoding. A fundamental limitation of this approach is the use of a predetermined bit size to encode phrases with varying lengths. Analysis of the deep semantic communication (deep-SC) scheme’s complexity, as shown in [51], reveals its superior performance compared to existing SSCC schemes. In contrast, the semantic communication technique based on JSCC exhibits reduced processing time compared to the deep-SC system. The successful incorporation [51] and analysis [19] of deep learning and JSCC techniques provide a valuable motivation for the innovative advancement of semantic communication. The fundamental objective of this work is to analyze the process of semantic source coding in connection to primary information. The objective of this framework is to surpass the constraints of Shannon’s conventional information theory, therefore enabling a more all-encompassing methodology.
The primary goal of deep learning-based Joint Source-Channel Coding (DJSCC) [52] is to reliably attain high levels of performance, even in scenarios characterized by resource constraints and low Signal-to-Noise Ratios (SNRs). The reconstruction scheme involves the use of DJSCC, as outlined in the research by [53], which may be seen as an adapted version of an auto-encoder. Supplementary Noise Ratios (SNRs) are included in the intermediate part of the encoder, resulting in the generation of a modified version of the original auto-encoder. In contrast to traditional wireless transmission techniques, the DJSCC encoder [54] employs semantic signals, which are analog waveform sequences, instead of digital transmission. Consequently, it may operate with high efficiency even in challenging channel conditions while still achieving enough recovery. Nevertheless, the enhanced efficiency attained by modern DJSCC techniques is only evident via simulations in scholarly research, which function as a type of information dissemination via supplementary transmission for semantic networks [55]. Typically, these models make the assumption of flawless synchronization, precoding, antenna, and power amplifier conditions. Therefore, the study [56] focuses on a DJSCC platform that employs SDR technology. The evaluation of the system’s capability relies on two crucial factors, namely non-linear distortion and synchronization inconsistencies. Moreover, the authors of [57] propose an innovative method for semantic communication, drawing inspiration from the impressive adaptability shown by Vision Transformers (ViTs) in successfully addressing various issues associated with image disruptive behavior. The potential advantages outlined in [58] of incorporating Unequal Error Protection (UEP) into a semantic encoder/decoder, using the current JSCC system, are worth investigating. The objective of this approach is to preserve the relative importance of data focused on semantic tasks. Furthermore, the implementation of a JSCC system may be simplified by using Quasi-cyclic Low-Density Parity-Check (QC-LDPC) codes on an adaptable device. A prototype for semantic communications is developed by constructing a fixed-point system that is then used to transmit and receive semantic information across a channel that contains residual noise.
In a recent publication, Zhang presented a new training approach for semantic communication systems. This approach builds upon the research accomplished by [59] and demonstrates its superiority compared to current digital and analog JSCC methods. A key challenge is in including modulation, a crucial step for effective transmission, into the training process. In order to address this problem, the authors use a theoretical framework that views modulation as a technique for data compression and employs noise to simulate its effects. This methodology enables the teaching of real-life situations and much exceeds current approaches in terms of efficacy. Two innovative approaches are proposed in the research of [60] to facilitate the seamless integration of digital communication and DJSCC. The first modulation, uniform modulation, functions as a flexible spring by modifying the spacing between data points (constellations) during transmission to align more effectively with the unique properties of the picture data being sent. The second technique, known as non-uniform modulation, adopts a more customized approach. The system examines the picture data and generates constellations which are particularly tailored to the distinct distribution of that particular image, therefore guaranteeing maximum transmission efficiency. Both methodologies ingeniously tackle the transformation of continuous analog data into discrete digital signals by meticulous adjustment of the system. The findings are remarkable, as this approach far surpasses current ways, thus facilitating enhanced picture transfer in digital communication. The work described in [61] introduces a novel approach called D2-JSCC to address the discrepancy in digital sources and thereby unlock the full potential of DJSCC. D2-JSCC operates as a tag team, where one component of digital source coding analyzes the image and identifies its semantic properties, taking into account the distribution of this statistical information.
In Fig.2, a semantic transmitter extracts and encodes meaningful information from the source, which is then modulated using light-emitting diode (LED)-based optical transmission [62]. The optical channel carries this data to a camera receiver, where it is demodulated and decoded through JSCC to recover the original semantic information. A key feature of the system is its use of feedback loops and knowledge sharing, which allow the transmitter to adapt its encoding and transmission strategies in response to channel conditions, improving robustness and efficiency. The system employs adaptive optical transmission, adjusting parameters like modulation and power levels to enhance performance under varying environmental conditions. The knowledge base in this semantic communication system plays a central role in enhancing efficiency and accuracy by focusing on transmitting meaningful information rather than raw data. The transmitter and receiver can interpret semantic representations by sharing a common knowledge base, ensuring efficient communication with reduced data transmission. The system incorporates real-time feedback, allowing the transmitter to adapt its encoding and transmission strategies based on current channel conditions, such as noise or interference.
Incorporating an LLM enhances the system’s ability to understand and generate contextually relevant content, improving the semantic accuracy of the transmitted information. Real-time feedback mechanisms allow the transmitter to adapt its encoding and transmission strategies based on current channel conditions, such as noise or interference. This adaptive approach optimizes LED-based optical transmission and employs joint source-channel coding, focusing on high-level semantic features instead of relying solely on traditional error correction methods. The knowledge base, augmented by the LLM, enables the system to intelligently manage noise and reconstruct lost data, resulting in a more bandwidth-efficient and power-optimized communication process, especially in dynamic environments. By leveraging the semantic understanding provided by the LLM, the system can deliver robust communication that aligns closely with user intent and context.
IV Enhancing Domain-Specific Semantic Effectiveness with Large Language Models
The development of LLMs has sparked significant interest in their ability to adapt to various domains. However, challenges remain, particularly in the areas of cognitive interference and memory retention when models are customized for specific tasks. Furthermore, the quest for a versatile LLM that can operate across multiple domains often leads to performance declines due to communication issues. To enhance LLM capabilities, integrating data fusion techniques becomes essential. Data fusion enables the synthesis of information from multiple sources, leading to richer context and improved decision-making. This integration can significantly enhance the model’s understanding and responsiveness in complex environments, such as networking and Earth observation. This discussion explores recent strategies and frameworks aimed at optimizing LLM performance, highlighting innovative approaches to enhance their adaptability, efficiency, and semantic processing through effective data fusion.
IV-1 Domain-adapted large language model
LLMs have difficulties in gaining various talents without encountering cognitive interference or memory deterioration when customized for particular domains. Furthermore, the endeavor to construct a versatile model capable of handling many domains simultaneously often results in a decline in overall performance due to the presence of miscommunication across the various domains. REGA presented in [63] has three strategies to tackle this issue: firstly, it uses illustrative examples to refresh the LLM’s memory of general information; secondly, it provides particular instructions for each new skill via customized prompts; and thirdly, it seamlessly incorporates a limited quantity of task-specific material with the existing general knowledge. However, creating customized deep-learning models for individual networking tasks is a laborious process that also faces difficulties when encountering novel scenarios. The authors of [64] propose using extensive LLMs that have been pre-trained on massive datasets. The NetLLM method, which follows a ”one model for all” strategy, seeks to address diverse networking issues by enhancing performance and adapting to new circumstances. Another factor to be taken into account is that a scarcity of labeled data impedes the training of text categorization algorithms. Domain Adaptation Large Language Model interpolator (DALLMi) presented in [65] employs a novel approach by using a small quantity of labeled data with a substantial volume of unlabeled data. It connects the two via word embeddings and a method known as interpolation. DALLMi additionally adjusts the training process to compensate for the uneven distribution of data. This strategy demonstrates superior performance compared to conventional approaches, particularly in situations when there is a limited availability of labeled data.
IV-2 Design of semantic effective plane
Although there has been notable progress in the development of LLMs, the field of semantic processing, which focuses on comprehending the meaning in language, seems to be experiencing a lack of significant innovation. By processing natural language input, it understands context, identifies key information, and generates coherent, contextually appropriate actions or solutions. For example, when asked to summarize a document, it doesn’t just extract random sentences—it analyzes the entire content, synthesizes the core ideas, and provides a concise summary. Similarly, when instructed to provide a recommendation, it evaluates the relevant data, processes the underlying patterns, and offers an informed suggestion. The key strength of a language model lies in its ability to turn complex text into clear, actionable outcomes, bridging the gap between human language and practical tasks.
OpenAGI [66] is an open-source AGI research platform designed to solve complex, multi-step tasks by synthesizing diverse models and utilizing datasets from sources like Hugging Face, GitHub, and LangChain. It introduces the Reinforcement Learning from Task Feedback (RLTF) approach, where a Large Language Model acts as a controller to select and execute external expert models, using task feedback to refine its planning strategy. By combining LLMs with domain-adapted models, OpenAGI aims to enhance task-solving capabilities, even with smaller-scale models, advancing the potential for Artificial General Intelligence (AGI). This is achieved by analyzing existing studies on these tasks and their practical applications, as illustrated in Fig.3 which depicts the semantic effective plane of Fig.1. Our proposed semantic effective plane can be achieved by the following outlines:
Figure 3: The proposed semantic effectiveness plane applying semantic processing with LLM technique
  • By leveraging knowledge graphs, OpenAGI enriches its understanding of complex tasks through a structured representation of entities and their relationships, allowing for more accurate and context-aware decision-making. This integration facilitates a deeper comprehension of the domain-specific knowledge and connections relevant to multi-step tasks.
  • In conjunction with knowledge graphs, OpenAGI utilizes pretrained predictive models to forecast potential outcomes and suggest optimal actions based on historical data and learned patterns. These models, trained on extensive datasets, provide valuable predictions that guide the LLM in selecting and synthesizing the most appropriate expert models for each task.
  • The semantic analyzer within OpenAGI extracts semantic importance from natural language queries and task descriptions. This capability enhances the platform’s ability to discern and prioritize critical elements within the text, ensuring that the most relevant information drives the task-solving process. By combining these components, OpenAGI not only synthesizes diverse models but also benefits from a deeper contextual understanding and predictive accuracy, advancing its potential for achieving AGI.
  • The hybrid mode of knowledge graphs and pretrained predictive models empowers OpenAGI to approach problems with a well-rounded perspective, integrating domain knowledge with predictive insights. As a result, OpenAGI not only refines its task-solving strategies through RLTF but also achieves a higher level of accuracy and efficiency in addressing complex, real-world challenges. OpenAGI can formulate task-oriented semantic commands. These commands are designed to be actionable and directly relevant to the task, guiding the execution of appropriate models or tools.
In modern networking architectures, different network layers serve distinct purposes, from physical transmission to data presentation, each playing a critical role in ensuring efficient communication. The semantic importance of these layers lies in their ability to abstract and compartmentalize network functions, such as error correction, flow control, routing, and data encryption. However, cross-layer inference, where decisions at one layer affect or leak into others, can introduce inefficiencies, security vulnerabilities, or performance bottlenecks. LLMs can mitigate these issues by understanding the dependencies between layers and applying intelligent decision-making strategies. By analyzing traffic patterns, latency, and data flow across layers, LLMs can propose optimizations that enhance end-to-end performance while maintaining layer independence.
IV-3 Automating Earth observation with LLMs integration
Domain-adapted LLM is promising to transform Earth observation by automating and improving analysis processes, making results more accessible to decision-makers. The framework presented in [67] leverages LLMs to interpret and orchestrate tasks using tools such as the Open Data Cube and OEA Algorithm Hub, effectively streamlining EO analysis and showcasing the potential of LLMs in enhancing geospatial data processing and decision support. GeoLLM-Engine harnesses the power of natural language to execute complex EO tasks by integrating geospatial API tools, dynamic maps, and multimodal knowledge bases [68]. Unlike existing systems that rely on simplified tasks and template-based prompts, GeoLLM-Engine evaluates agent performance on realistic, high-level commands through a massively parallel engine using 100 GPT-4-Turbo nodes, handling over half a million multi-tool tasks and 1.1 million satellite images. This approach advances beyond traditional single-task methods to assess state-of-the-art agents and prompting techniques with long-horizon prompts.
V Semantic feature processing for Resource-efficient communication and interference cancellation
This section introduces a Cognitive Semantic EO system model designed to enhance satellite communication networks by integrating advanced semantic processing techniques, such as Deep Joint Source-Channel Coding (DT-JSCC) and Semantic Data Augmentation (SA). The model leverages inter-satellite links and cognitive semantic augmentation to optimize the transmission and analysis of multispectral EO data. By focusing on transmitting essential, task-relevant semantic features rather than raw data, the system reduces communication overhead and energy consumption, while improving decision-making and performance in 6G satellite communication networks. The proposed framework showcases an end-to-end semantic-augmented communication process, ensuring efficient data handling and enhanced user experience.
V-A Cognitive Semantic EO system model and problem formulation
Figure 4: The proposed semantic cognitive satellite network applying DT-JSCC and semantic data augmentation technique.
In this satellite system illustrated in Fig.4, a low Earth orbit (LEO) satellite establishes communication with a gateway located at a ground terminal directly connected to a server. Moreover, the satellite is smoothly integrated into an intricate network of inter-satellite links, allowing it to exchange real-time data with other spacecraft in orbit. Implementing an inter-satellite connection improves data transmission reliability and extends the communication period before the satellite moves outside the ground terminal’s coverage area.
The system processes multispectral satellite images captured at different time points, utilizing two satellites. The Sat 1, positioned near the ground terminal, captures the reference image at time $t_{0}$, denoted as $U(t_{0})$. This image is represented as $U(t_{0})=\{u^{t_{0}}(i,j,k)\mid 1\leq i\leq H,1\leq j\leq W,1\leq k\leq D\}$, where $H$ is the height, $W$ is the width, and $D$ is the number of spectral bands.
V-A1 Semantic data transmission
Upon cognitive processing of the reference image $U(t_{0})$, semantic information $U_{sem}(t_{0})$ is generated by the classification network $\textit{f}_{s1}$ of on-board processor on Sat 1. This semantic information can be mathematically represented as $U_{sem}(t_{0})=\textit{f}_{s1}(U(t_{0}))$, where $U_{sem}(t_{0})=\{u^{t_{0}}(i,j,k)\mid 1\leq i\leq B,1\leq j\leq C,1\leq k\leq A\}$. We define $B$ as the batch size, $C$ as the number of classes, and $A$ as the feature dimension. This information is then sent across an inter-satellite communication connection to the second satellite, which is located at a further distance and serves as a relay. The transmission and receiving procedure can be executed using the DT-JSCC approach [69], which relies on a certain quantization level to provide an error correction for channel coding. As illustrated in Fig.5 at time $t_{1}$, Sat 2 employs the semantic information $U_{sem}(t_{0})$ transmitted by Sat 1 encoding $\textit{f}_{s2}$ to examine and establish a correlation to the current image $U(t_{1})$ that Sat 2 has obtained. This operation can be achieved by on-board training for more semantic features. This ensures that Sat 2 can effectively analyze the new image in the context of the reference image received from Sat 1.
V-A2 Cognitive semantic augmentation
Sat 2 utilizes semantic information $U_{sem}(t_{0})$ to extract and acquire pertinent features from the $U(t_{1})$ image. Semantic cognitive augmentation intends to prioritize both the pertinent data for the task and the semantic feature invariant, facilitating the effective detection of important changes and correlations between the reference and new images through the Semantic Data Augmentation (SA) methodology [70]. As illustrated in Fig.5, the covariance matrix prediction network $\textit{g}_{s2}$ learns from $U_{sem}(t_{0})$ to minimize the SA training loss $\textit{L}_{SA}(\textit{f}_{s2};\textit{g}_{s2}|U_{sem}(t_{0}))$. This can be done by setting up an SA task and optimizing $\textit{g}_{s2}$ by backpropagating the SA-gradient $\theta_{\textit{g}_{s2}}$ and minimizing the cross-entropy loss of the output of DT-JSCC decoder $\textit{l}_{s2}$. This technique aligns with the objective of the system to minimize both data transmission and on-board energy consumption [71]. Processing data on the semantic feature domain $U_{sem}(t)$, as opposed to the multi-spectral pixels domain $MP(t)$, shows potential for transmitting cognitive information effectively to satellites in close proximity. We provide the semantic cognitive covariance matrix $\upsilon_{t_{1}}^{\textit{g}_{s2}}=\textit{g}_{s2}(U_{sem}(t_{0}),\theta_{\textit{g}_{s2}})$, which enhances the generalization of image change in EO data and facilitates a smooth integration into the optimization objective of the classification network. The scope of this extends beyond particular tasks linked to the development of certain neural networks. Therefore, the cognitive semantic channel can be facilitated by the above methodology of intergrating DT-JSCC and semantic augmentation technique, which expands the applications of classification tasks for 6G communication.
Definition 1.
Cognitive semantic augmentation is the enhancement of a cognitive agent’s ability to interpret and process information by leveraging semantic knowledge. It enables the agent to prioritize relevant features, maintain semantic consistency, and improve the detection of important changes and patterns, thereby enhancing decision-making and understanding in complex environments.
Figure 5: The proposed transmitter and receiver for cognitive semantic satellite networks.
V-A3 End-to-end semantic-augmented communication
As illustrated in Fig. 4, the end-to-end communication between EO Sat 1, EO Sat 2, and the User Terminal (UT) involves several advanced steps to optimize the transmission and interpretation of EO data using the concept of Cognitive Semantic Augmentation (CSA). EO Sat 1 initiates the process by collecting signals from the Earth’s surface, which could be in the form of imagery, radar, or other sensor data. Rather than transmitting raw data, Sat 1 applies DT-JSCC to compress the data into quantized semantic features, retaining only the most essential information needed for specific tasks, thus reducing communication overhead. These semantic features are transmitted over an inter-satellite link to EO Sat 2, which is enhanced with semantic processing capabilities. Upon receiving the data, EO Sat 2 conducts on-board semantic inference and augmentation processing, analyzing and further interpreting the features to extract meaningful insights, such as detecting objects, patterns, or changes over time. The processed semantic data is then transmitted to the User Terminal (UT), which allows the UT to access this refined information. Before EO Sat 2 transmits its downlink data, UT identically received the Sat1 EO data $U_{sem}(t_{0})$ to perform SA technique [70] and optimize DT-JSCC decoder $\textit{l}_{UT}$. Therefore, the classification networks $\textit{f}_{s2}$ and $\textit{l}_{UT}$ can be updated individually as semantic-enhanced DT-JSCC to facilitate the CSA downlink. As illustrated in Fig.5, we summarize the proposed end-to-end CSA satellite communication in Algorithm 1.
Algorithm 1 End-to-end CSA Satellite communication
Input: Sat1 and Sat2 EO image $U(t_{i})$: Height $H$, Width $W$, and $D$ spectral bands; Semantic information: batch size $B$, $C$, number of classes $C$, and feature dimension $A$;
Output: UT EO inference;
1:for $i=0,1,\cdots,n-1$ do
2:   Sat1 processing and transmit to Sat2 and UT: $U_{sem}(t_{i})=\textit{f}_{s1}(U(t_{i}))$.
3:  Sat2 receive $U_{sem}(t_{i})$ and processing $U_{sem}(t_{i+1})=\textit{f}_{s2}(U(t_{i+1}))$.
4:  Covariance prediction network of Sat2: Meta-learning($\textit{g}_{s2}$,$\textit{f}_{s2}$,$\textit{l}_{s2}$) to process:
5:     $\upsilon_{t_{i+1}}^{\textit{g}_{s2}}\rightarrow$Minimise $\textit{L}_{SA}(\textit{f}_{s2};\textit{g}_{s2}|U_{sem}(t_{i}))$ and Optimizing $\textit{g}_{s2}$ by backpropagating $\theta_{\textit{g}_{s2}}$.
6:   Covariance prediction network of UT: Meta-learning($\textit{g}_{UT}$,$\textit{f}_{UT}$,$\textit{l}_{UT}$) to process:
7:   $\upsilon_{t_{i+1}}^{\textit{g}_{UT}}\rightarrow$Minimise $\textit{L}_{SA}(\textit{f}_{UT};\textit{g}_{UT}|U_{sem}(t_{i}))$ and Optimizing $\textit{g}_{UT}$ by backpropagating $\theta_{\textit{g}_{UT}}$.
8:   Sat2 transmit SA $U_{sem}(t_{i+1})$ to UT $\textit{l}_{UT}$ for SA EO inference.
9:end for
Once the semantic inference is complete, the data undergoes semantic reconstruction, processing, and analysis to produce actionable insights, which are then made available to user applications through Semantic Feature Division Access (SFDMA) [72]. These user applications can use the semantically processed EO data for a variety of purposes, including inter-UE (User Equipment) interference cancellation, which improves communication efficiency and signal clarity. The SFDMA network reduces interference through approximate orthogonal transmission and introduces the Alpha-Beta-Gamma formula for semantic communications, with simulations validating its effectiveness. This overall process demonstrates how integrating semantic cognition into satellite systems can drastically improve the efficiency of EO data transmission, interpretation, and utilization, minimizing unnecessary data loads and focusing on task-relevant information.
V-A4 On-board processing efficiency
The authors presented in [73] provide a detailed comparison of several deep learning models, focusing on metrics like input size, total parameters, estimated model size, FLOPs, and the time taken for training and inference. Simpler architectures, such as CNN and ResNet-50, exhibit low computational complexity. For instance, CNN has the smallest total parameters (0.07 MB) and FLOPs (0.93 MFLOPs), making it highly efficient for training (233 seconds) and inference (7 ms). ResNet-50, while more complex, still maintains moderate efficiency, with 195.7 million parameters, a model size of 100 MB, and 11.78 MFLOPs, resulting in faster inference times (0.83 s) compared to more complex models.
Transformer-based models, such as ViT, SwinTransformer, and EfficientNet variants, are far more computationally intensive. ViT-large (Google) leads the complexity spectrum with 303 million parameters, 162.20 GFLOPs, and significant memory requirements (1.24 GB), translating into much longer training (9,271 s) and inference (2 s) times. SwinTransformer and EfficientNet-V2 models offer a compromise between complexity and efficiency. SwinTransformer, for instance, has 85.7 million parameters, 21.57 GFLOPs, and takes around 538 seconds for training, making it efficient for high-performance tasks despite its complexity. The authors highlights that transformer models like ViT-huge (Google) and ViT-large are the most resource-intensive, while smaller models like SmallViT and EfficientNet-V2 maintain a good balance between model size, computational complexity, and time efficiency. Ultimately, this comparison shows a trade-off between performance and computational demand, with traditional CNN models being more efficient for simpler tasks, and transformer models excelling at complex, large-scale tasks but requiring significantly more computational resources.
V-B Numerical Results
V-B1 Channel model
Let $X$ denote the transmitted signal and $Y$ denote the received signal. The general form of the channel equation is:$Y=H\cdot X+N$, where $H$ is the channel gain and $N$ is the noise term. Below are the specific formulations for different types of channels.
For an AWGN channel, there is no fading, so the channel gain $H=1$: $Y=X+N_{\text{AWGN}}$, where $N_{\text{AWGN}}\sim\mathcal{N}(0,\sigma^{2})$ is the Gaussian noise with variance $\sigma^{2}$. In a Rician channel, the channel gains $H$ follows a Rician distribution: $Y=H_{\text{Rician}}\cdot X+N$, where $H_{\text{Rician}}$ is a Rician-distributed random variable characterized by the Rician $K_{r}$factor, and $N\sim\mathcal{N}(0,\sigma^{2})$ is the noise. For a Rayleigh channel, the channel gain $H$ follows a Rayleigh distribution: $Y=H_{\text{Rayleigh}}\cdot X+N$, where $H_{\text{Rayleigh}}$ is a Rayleigh-distributed random variable, and $N\sim\mathcal{N}(0,\sigma^{2})$ is the noise.
In the LEO Rician channel, the received signal incorporates the Rician fading characteristics along with effects like Doppler shift and satellite-specific parameters: $Y=H_{\text{LEO-Rician}}\cdot X+N_{\text{AWGN}}$, where $H_{\text{LEO-Rician}}$ is a Rician-distributed fading gain in the LEO environment and includes noise contributions such as scintillation and atmospheric loss according to Table II. For a LEO Rayleigh channel, the fading gain follows a Rayleigh distribution with LEO-specific noise characteristics: $Y=H_{\text{LEO-Rayleigh}}\cdot X+N_{\text{AWGN}}$, where $H_{\text{LEO-Rayleigh}}$ is a Rayleigh-distributed fading gain.
To simulate propagation and attenuation, the channel model incorporates both small- and large-scale fading; the path loss components are the large-scale fading between the satellite and UT [74]: $P_{L_{\text{tot}}}=P_{L_{b}}+P_{L_{g}}+P_{L_{s}}$, in which the attenuation resulting from atmospheric gases is denoted by $P_{L_{g}}$, the attenuation resulting from ionospheric or tropospheric scintillation by $P_{L_{s}}$, and the basic path loss is represented by $P_{L_{b}}$. Furthermore, the ISL of free space may be understood as $P_{L_{\text{ISL}}}=P_{L_{ib}}$, in which $P_{L_{ib}}$ denotes Sat1 and Sat2’s fundamental path loss. These route loss components are all expressed in decibels (dB). More specifically, the free space propagation and shadow fading of the signal are taken into account using the fundamental path loss model $P_{L_{b}}$. For a distance $d_{k,i}$ also known as slant range, the free space path loss ($P_{L_{\text{FS}}}$) in dB range in meters and frequency $f_{c}$ in GHz. Hence, the slant distance $d_{k,i}$ can be expressed as [75]:
$d=\sqrt{R_{E}^{2}\sin^{2}\theta+r_{m}^{2}+2R_{E}r_{m}-2R_{E}r_{m}\sin\theta},$ (1)
where $R_{E}=6378\,\text{km}$ is the radius of the Earth, $r_{m}$ and $\theta$ are the satellite altitude and its elevation angle, respectively.
Thereby, the free space path loss ($P_{L_{\text{FS}}}$) can be calculated as: $P_{L_{\text{FS}}}(d,f_{c})=32.45+20\log_{10}(f_{c})+20\log_{10}(d)$, Shadow fading (SF) is modeled by a log-normal distribution as $\mathcal{N}(0,\sigma^{2}_{SF})$ with zero-mean and $\sigma_{SF}$ standard deviation. The values of $\sigma^{2}_{SF}$ can be extracted from the 3GPP Release-15. Then, the path loss with and without shadow fading in dB units is modeled as $P_{L_{b}}=P_{L_{\text{FS}}}(d,f_{c})+SF$ and $P_{L_{ib}}=P_{L_{\text{FS}}}(d_{ib},f_{c})$, where $d_{ib}$ represent the distance of Sat1 and Sat2.
We consider the Rician model to express the channel between LEO satellite $s1$ and the UT $z$ and ISL between LEO satellite $s1$ and $s2$. The channel from the LEO satellite $s1$ to the UT $z$ and ISL are given as
$f_{s1,z}=\sqrt{\frac{R_{z}\zeta_{s1,z}}{R_{z}+1}}\bar{f}_{s1,z}+\sqrt{\frac{\zeta_{s1,z}}{R_{z}+1}}\tilde{f}_{s1,z},$ (2)
$f_{s1,s2}=\sqrt{\frac{R_{s2}\zeta_{s1,s2}}{R_{s2}+1}}\bar{f}_{s1,s2},$ (3)
where $\bar{f}_{s1,z}$, $\bar{f}_{s1,s2}$ and $\tilde{f}_{s1,z}$ are the line-of-sight (LoS) and non-line-of-sight (NLoS) paths components, respectively, $R_{z}$ is the Rician factor, and the ISL does not have NLoS component. After accounting for the LEO satellite antenna gain $G_{T}$, the large-scale fading effects of the LEO satellite $s1$ to the UT $z$ and ISL may be modeled as follows: $\zeta_{s1,z}(\text{dB})=PL_{\text{tot}}-G_{T}$ and $\zeta_{s1,s2}(\text{dB})=PL_{\text{ISL}}-G_{T}$. Furthermore, the complex Gaussian distribution with zero mean and unit variance is followed by the independent and identically distributed (i.i.d.) entries of $\tilde{f}_{s1,z}$ data.
The complex baseband downlink space domain channel at a time instance $t$ and frequency $f$ can be expressed as follows using a ray-tracing-based channel modeling method:
$H_{s1,z}(t,f)=f_{s1,z}\exp\left(2\pi\left[tv_{z}-f\tau_{z}\right]\right),$ (4)
$H_{s1,s2}(t,f)=f_{s1,s2}\exp\left(2\pi\left[tv_{s2}-f\tau_{s2}\right]\right),$ (5)
where $v_{z}$, $v_{s2}$ are the Doppler shift, and $\tau_{z}$, $\tau_{s2}$ are the propagation delay. Table II summarises the simulation parameters.
TABLE II: Simulation Parameters
Parameter Value
Carrier frequency 28 GHz
Satellite antenna gain 35 dBi
User antenna gain 37 dBi
Scintillation loss 0.5 dB
Atmospheric loss 0.3 dB
Rician factor 2.8
V-B2 Physical layer transmission
Fig. 6 and Fig. 7 present Top-1 accuracy using DT-JSCC under different channel conditions for EuroSAT data. The top two graphs depict performance across different PSNR (Peak Signal-to-Noise Ratio) levels for the 16PSK modulation scheme. We show accuracy for the AWGN (Additive White Gaussian Noise) and Rician channels, with different K-factor values (32, 64, 128), where a higher K-factor generally corresponds to better channel quality. Furthermore, we compare accuracy across the Rician and Rayleigh channels for the same K values, showing higher accuracy in the Rician channel. The lower two graphs show similar comparisons but for the LEO environment with 16APSK modulation. We examine the AWGN and Rician channels, while the right-hand side evaluates Rician and Rayleigh channels under different K-factors. Across all scenarios, the graphs indicate that increasing PSNR improves accuracy, with Rician channels generally performing better than Rayleigh, and higher K-factors leading to more stable performance, especially in LEO environments.
Figure 6: Top1 accuracy using DT-JSCC based on 16PSK/16APSK AWGN and Rician channel.
Figure 7: Top1 accuracy using DT-JSCC based on 16PSK/16APSK LEO Rician and LEO Rayleigh channel.
In comparing the confusion matrices of Fig.8, we can observe some key performance differences over the AWGN channel where PSNR=14dB. For k=128, the classification accuracy across most categories is notably high, with certain categories like ”Forest” showing a strong prediction rate of 99.21%, while there are minor misclassifications in categories like ”PermanentCrop” and ”Pasture” with accuracies of 93.71% and 94.59% respectively. Meanwhile, the case of K=32 also performs well but demonstrates slightly different classification behaviors, such as a stronger prediction for ”River” at 98.05% but more substantial confusion in categories like ”Forest” and ”Residential,” where there are larger misclassification rates with accuracies of 95.44% and 90.88% compared to the K=128 performance. The identical trend of K=32 and K=128 can be observed over the 16APSK Rician channel. However, Fig.9 demonstrates better accuracy and classification than 16PSK modulation in almost all categories, while there are minor misclassifications in categories like ”HerbaceousVegetation” and ”PermanentCrop” with accuracies of 98% and 98.34% respectively. It is noteworthy that 16APSK modulation validated where PSNR at 12dB is better than 16PSK at 14 dB.
Figure 8: Top1 accuracy of confusion matrix using DT-JSCC based on 16PSK AWGN channel where PSNR=14dB (Upper K=128 and lower K=32).
Figure 9: Top1 accuracy of confusion matrix using DTJSCC based on 16APSK Rician channel where PSNR=12dB and K=128.
V-B3 CSA satellite network
Table III shows CSA DT-JSCC system with high accuracy and evidently with most classes, such as Annual Crop (94.51%), Herbaceous Vegetation (91.43%), and Sea/Lake (98.85%), showing high Top1 accuracy, indicating correct classifications. Some misclassifications can be improved, such as between Industrial and Residential areas. The non-CSA shows slightly reduced accuracy, with some increased misclassifications, particularly in Herbaceous Vegetation (89.64%) and Sea/Lake (96.63%). Overall, the CSA-enhanced system demonstrates better accuracy and fewer misclassifications compared to the non-CSA system.
TABLE III: Top1 accuracy comparison of CSA and non-CSA DT-JSCC K=32 systems while PSNR=12dB and 16APSK over LEO Rician channel.
Class Accuracy (%) CSA Accuracy (%) non-CSA
AnnualCrop 94.50 88.38
Forest 98.41 98.84
HerbaceousVegetation 91.63 89.64
Highway 92.30 77.27
Industrial 95.95 90.95
Pasture 94.59 90.53
PermanentCrop 90.03 83.55
Residential 98.02 95.95
River 92.33 72.63
SeaLake 98.95 98.33
As illustrated in Fig. 10, the SA technique struggles significantly under the conditions of a low PSNR value due to the inherent noise and interference present in LEO Rician channels. This adverse environment leads to diminished signal quality, which adversely impacts the performance of the SA method. In contrast, a notable enhancement in Top1 accuracy is evident with both K=32 and K=128 DT-JSCC transmissions. This improvement can be attributed to the end-to-end CSA architecture, which leverages advanced techniques to optimize signal processing, resulting in a more robust transmission. By intelligently aggregating information, the system can better maintain signal integrity, thereby improving overall accuracy in challenging channel conditions.
Figure 10: Top1 accuracy of CSA satellite network using DT-JSCC over 16APSK LEO Rician channel.
The results highlight the importance of employing sophisticated transmission strategies, particularly in environments characterized by high levels of noise and interference. The substantial gains in accuracy achieved with the DT-JSCC approach underscore the efficacy of incorporating channel-aware mechanisms in the design of communication systems, suggesting that further exploration and development in this area could yield even more significant benefits in future applications.
VI Conclusion
In this paper, we explored the critical role of semantic inference and deep learning in enhancing EO systems, focusing on the integration of data-driven modeling approaches. With the rise of big EO data and the increasing complexity of satellite imagery, conventional methods struggle to meet the demands of modern applications such as environmental monitoring, disaster management, and resource allocation. By utilizing advanced deep learning techniques, we have illustrated the potential to overcome key challenges, including the scarcity of labeled data, the variability in data sources, and the intricacies of multi-dimensional datasets. The incorporation of semantic inference enables systems to extract meaningful insights from vast and complex datasets, paving the way for more accurate decision-making in both civilian and critical operations.
This work presents an advanced framework for semantic communication networks, with a focus on the development and optimization of Cognitive Semantic EO systems. By incorporating semantic-aware methodologies like JSCC and DT-JSCC, the proposed system enhances communication efficiency, particularly in satellite networks. The integration of semantic data augmentation, inter-satellite links, and cognitive processing capabilities ensures that only the most relevant, task-specific data is transmitted, significantly reducing communication overhead and improving system performance. Furthermore, by leveraging semantic inference and cognitive augmentation techniques, the system facilitates better image analysis, change detection, and decision-making in real-time applications. The proposed end-to-end model demonstrates how integrating semantic cognition into satellite systems can drastically improve data transmission, interpretation, and overall system efficiency, thereby addressing the demands of 6G networks and beyond. This research lays the groundwork for future advancements in semantic communication, with promising applications in a wide range of fields, including remote sensing, autonomous systems, and smart communication networks.
Acknowledgment
This work was funded by the Luxembourg National Research Fund (FNR), with the granted SENTRY project corresponding to grant reference C23/IS/18073708/SENTRY.